OpenStudy (anonymous):

Does this limit exist, if so how would I solve it, lim x->3 (1/x)((1/(x+3)^2)-(1/9))

OpenStudy (imstuck):

Is that this:\[\lim_{x \rightarrow 3}(\frac{ 1 }{ x })(\frac{ 1 }{ (x+3)^{2} }-\frac{ 1 }{ 9 })\]

OpenStudy (anonymous):

Yes

OpenStudy (imstuck):

That's it exactly?

OpenStudy (anonymous):

Yep

OpenStudy (imstuck):

This can be solved by direct substitution of 3 in for x. There is no vertical asymptote created when you insert a 3. Try plugging in 3 for x. I did and got an answer of -1/36

OpenStudy (anonymous):

Okay thats what I had but I thought that would be wrong, mind if I ask another question?